OpenStudy (anonymous):

Simplify quantity 6 times x plus 24 all over 6

OpenStudy (anonymous):



OpenStudy (anonymous):

6x+24 ---- 6

OpenStudy (anonymous):

x+4 right? just divide each term by 6

OpenStudy (anonymous):

If it makes it easier for you to see, you can break that into two fractions! \[\large \frac{6x+24}{6} = \frac{6x}{6} + \frac{24}{6}\]

OpenStudy (anonymous):

thankss again! :D

OpenStudy (anonymous):

My pleasure! =D